S.T a rectangle of maximum perimeter which can be inscribed in a circle of radius r is a square of side root2*r? 

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Linta

let the side of rectangle is a and b and its center coinside with the center of circle.

a/2 = r cosΘ  then  b/2 =r sinΘ


now perimeter P =2a+2b

                        =4r(cosΘ + sinΘ)

                         = 4r√2sin(Θ +45)

so for maximum P   , sin(Θ +45) =1

              so Θ =45

  so a=b = r√2

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin